OpenStudy (anonymous):

A snail was observed to cover 22 cm in 20 seconds and then another 21 cm in 30 seconds. What was the average speed of the snail? A. 0.7cm/s B. 0.86cm/s C. 0.9cm/s D. 1.1cm/s

OpenStudy (anonymous):

use S=VT here S=distance coverd=22+21=43cm V=velocity=? T=time=20+30=50sec so S=VT =>V=S/T=43/50=0.86cm/sec